# RLAIF: Scaling Reinforcement Learning from Human Feedback with AI Feedback

### v2

@@ 3.4 Human Evaluation
For experiments evaluated by win rates, evaluators were presented with an input context and multiple responses generated from different policies (e.g. RLAIF, RLHF, and SFT). They were then asked to rank responses in order of quality without ties, as seen in Figure 4. Input contexts were drawn from test splits of datasets, which were not used for training or any other evaluation [^6] For summarization, we used the test split of Reddit TL;DR. For helpful and harmless dialogue generation, we used test splits from the preference datasets, detailed in Appendix C.. Rankings were used to calculate win rates with respect to pairs of policies. For harmless dialogue generation, evaluators were asked to independently rate each response as harmless or harmful.
@@ 4 Results
Direct RLAIF vs SFT 74%
Direct RLAIF vs Same-size RLAIF 60%
Table 1: Left side: Win rates when comparing generations from two different models for the summarization and the helpful dialogue tasks, judged by human evaluators. Right side: Harmless rates across policies for the harmless dialogue task, judged by human evaluators.
@@
RLAIF achieves performance gains on par with or better than RLHF on all three tasks (see Figure 1 and Table 1). RLAIF and RLHF are preferred by human evaluators over the baseline SFT policy 71% and 73% of the time for summarization [^7] RLAIF and RLHF are also preferred over the human reference summaries in Reddit TL;DR 79% and 80% of the time, respectively. and 63% and 64% for helpful dialogue generation, respectively. The difference in win rates between RLAIF vs. SFT and RLHF vs. SFT are not statistically significant. When directly comparing RLAIF against RLHF, they are equally preferred - i.e. the win rate is not statistically significantly different from 50%. For harmless dialogue generation, RLAIF achieves a harmless rate of 88%, outperforming both RLHF and SFT, which score 76% and 64%, respectively [^8] RLAIF achieves a statistically significant improvement over RLHF and SFT, according to a two-sample t-test..
@@
As observed in Stiennon et al. (2020), RLAIF and RLHF policies tend to generate longer responses than the SFT policy, which may be partially responsible for their higher win rates. We conduct post-hoc analysis to control for length and find that both RLAIF and RLHF policies still outperform the SFT policy, and by similar margins to one another. See Appendix J for details.
@@
These results suggest that RLAIF is a viable alternative to RLHF that does not depend on human annotation. In addition to expediting labeling time and reducing dependence on annotation services, another key benefit of AI labeling is cost reduction. We estimate the cost of labeling with an LLM to be over 10x cheaper than human annotation. See Appendix L for detailed calculations.
@@ 4.2 Towards Self-Improvement
In Section 4.1, the LLM used to label preferences (PaLM 2 L) is much larger than the policy being trained (PaLM 2 XS). Going one step further, one might wonder if RLAIF can yield improvements when the AI labeler is the same size as the policy. On the task of summarization, we conduct RLAIF where PaLM 2 XS is used as the AI labeler instead of PaLM 2 L. The rest of the setup mimics the experiment in Section 4.1. We refer to this setup as “same-size RLAIF”.
Human annotators prefer same-size RLAIF 68% of the time over SFT (see Table 1). For reference, RLAIF using an AI labeler larger than the policy is preferred 71% over SFT [^9] The difference between win rates between “same-size RLAIF vs. SFT” and “RLAIF vs. SFT” is not statistically significant. For a two-sample t-test, p-value = 0.07. At alpha = 0.05, this difference is not statistically significant.. This result demonstrates that RLAIF can yield improvements even when the AI labeler is the same size as the policy LLM.
We note that the AI labeler and initial policy are not the exact same model. The AI labeler is the instruction-tuned PaLM 2 XS, whereas the initial policy is PaLM 2 XS fine-tuned on Reddit TL;DR summarization. Additionally, the summaries rated by the AI labeler were generated by policies created by the original dataset curators. For these reasons, we do not consider this experiment a strict case of “self-improvement”(Huang et al., 2022). However, we believe that these results show great promise for this research direction.
4.3 Direct RLAIF
In Sections 4.1 and 4.2, AI feedback was distilled into a RM. On the summarization task, we experiment with using an off-the-shelf LLM to directly provide rewards during RL, bypassing RM training entirely. Since using a large AI labeler in RL is computationally expensive, we use the smaller instruction-tuned PaLM 2 XS as the off-the-shelf LLM. We refer to this setup as “direct RLAIF”.
Human annotators prefer responses from direct RLAIF 74% of the time over SFT responses (see Table 1). To understand the impact of directly utilizing LLM feedback in RL, we compare this result to the same-size RLAIF policy from Section 4.2, which solely differs in training a RM that provides rewards during RL. Direct RLAIF outperforms same-size RLAIF, which achieves a statistically significantly lower win rate of 68%. Furthermore, when shown responses side-by-side, raters prefer direct RLAIF over same-size RLAIF 60% of the time [^10] This is statistically significantly different from 50% according to a two-sample t-test.. One hypothesis for the improved quality is that bypassing the distillation from AI preferences into a RM enables information to flow directly from the off-the-shelf LLM to the policy.
@@ 4.4 Prompting Techniques
Prompt
Summary
H1
H2
Base 0-shot
76.1%
67.8%
69.4%
Base 1-shot
76.0%
67.1%
71.7%
Base 2-shot
75.7%
66.8%
72.1%
Base + CoT 0-shot
77.5%
69.1%
70.6%
Detailed 0-shot
77.4%
67.6%
70.1%
Detailed 1-shot
76.2%
67.6%
71.5%
Detailed 2-shot
76.3%
67.3%
71.6%
Detailed 8-shot
69.8%
Detailed + CoT 0-shot
78.0%
67.8%
70.1%
Detailed + CoT 1-shot
77.4%
67.4%
69.9%
Detailed + CoT 2-shot
76.8%
67.4%
69.2%
Table 2: We observe that eliciting chain-of-thought reasoning tends to improve AI labeler alignment, while few-shot prompting and detailed preambles have mixed effects across tasks. H1 refers to helpfulness, H2 to harmlessness.
@@
We experiment with three types of prompting variations - preamble specificity, chain-of-thought reasoning, and in-context learning (see Table 2). We observe that eliciting chain-of-thought reasoning generally improves AI labeler alignment, while the impacts of preamble specificity and in-context learning vary across tasks. The best prompts outperform the base prompts (“Base 0-shot”) by +1.9%, +1.3%, and +1.7% for summarization, helpfulness, and harmlessness, respectively.
@@
Detailed preambles consistently improve alignment for summarization, while yielding mixed results for helpful and harmless dialogue generation. We hypothesize that summarization benefits more from a specific preamble due to the high complexity of this task. On the other hand, rating helpfulness and harmlessness are more intuitive to grasp, and therefore may benefit less from detailed instructions.
@@
Chain-of-thought reasoning improves alignment consistently for summarization. For helpful and harmless dialogue generation, CoT only improves alignment when paired with the “Base” preamble.
@@
Surprisingly, we observe that few-shot in-context learning only improves alignment for harmless dialogue generation [^11] We verified that all inputs used in these experiments fit within our AI labeler’s context length.. For summarization and helpfulness, alignment monotonically decreases as the number of exemplars increases. It seems unlikely that this effect is a result of exemplar quality, as exemplars were carefully handpicked to be high-quality and representative of each preference task. Furthermore, we conducted 10 trials for “Base 1-shot” on summarization, where a different exemplar was randomly selected for each trial. The maximum AI labeler alignment from all trials was 76.1%, which still did not surpass “Base 0-shot” in terms of AI labeler alignment. One hypothesis for why exemplars do not help is that the summarization and helpful dialogue generation tasks may already be sufficiently well-understood by the powerful AI labeler, rendering the exemplars unhelpful or distracting. It’s interesting to note that in-context learning is still an important research area that is not fully understood (Min et al., 2022; Wang et al., 2022a).
@@
For summarization, we compare against human inter-annotator agreement to get a sense of how well our LLM labeler performs in absolute terms. Stiennon et al. (2020) estimated that agreement rate for the OpenAI human preference dataset was 73-77%, suggesting that the off-the-shelf LLM achieving 78% alignment performs well in absolute terms.
@@
We also conduct experiments with self-consistency (Wang et al., 2022b), where multiple chain-of-thought rationales are sampled with temperature $T>0$. The preference distributions generated by the LLM are averaged together to arrive at the final preference label. We find that self-consistency strictly degrades AI labeler alignment (see Appendix M).
@@
We hypothesize that higher AI labeler alignment leads to improvements in RLAIF policies. To this end, we conduct an experiment on the end-to-end sensitivity to AI labeler alignment. Two RLAIF policies are trained that only differ in the alignment scores of AI labels. Results show that the policy trained with more aligned AI labels achieves a significantly higher win rate. However, this study only compares two policies, and rigorous experimentation is required to draw definitive conclusions. See Appendix N for details.
@@ 4.5 Size of LLM Labeler
Large model sizes are not widely accessible and can be slow and expensive to run. On the task of summarization, we experiment with labeling preferences with varying LLM sizes and observe a strong relationship between size and alignment (see Table 3). Alignment decreases -4% when moving from PaLM 2 Large (L) to PaLM 2 Small (S), and decreases another -11% when moving down to PaLM 2 XS - a trend consistent with scaling behaviors observed in other work (Kaplan et al., 2020). Besides general model capability, another contributing factor to this trend may be that smaller LLMs are more susceptible to position bias (see Appendix B).
On the other end of this trend, these results also suggest that scaling up AI labeler size may produce even higher quality preference labels. Since the AI labeler is only used to generate preference examples once and is not called during RL, using an even larger AI labeler is not necessarily prohibitively expensive.
Model Size
AI Labeler Alignment
@@
PaLM 2 S
73.8%
PaLM 2 XS
62.7%
@@ 5 Qualitative Observations
To better understand how RLAIF compares to RLHF, we inspected responses generated by both policies for the summarization task. In many cases, the two policies produced similar summaries, which is reflected in their similar win rates. However, we identified two patterns where they sometimes diverged.
The first pattern we observed is that in some cases, RLAIF hallucinates when RLHF does not. The hallucinations in RLHF summaries sound plausible but are inconsistent with the original text. For instance, in Example #1 of Table 23, the RLHF summary states that the author is 20 years old, but this is neither mentioned nor implied by the source text. The second pattern we observed is that RLAIF sometimes produces less coherent or grammatical summaries than RLHF. For instance, in Example #1 of Table 24, the RLAIF summary generates run-on sentences.
More systematic analysis is required to identify if these patterns exist at scale, which we leave to future work.
@@ 6 Related Work
LLMs have shown impressive performance over a wide range of NLP tasks (Brown et al., 2020; Thoppilan et al., 2022; Chowdhery et al., 2022; Google et al., 2023; OpenAI, 2023a). For several of these tasks, RL has emerged as an effective optimization technique. While initial applications of RL on tasks such as translation (Wu et al., 2016, 2018) and summarization (Gao et al., 2019; Wu and Hu, 2018) used automatic evaluation metrics as rewards, such simplified formulations of rewards did not fully align with human notions of quality.
Reinforcement learning from human feedback (Christiano et al., 2017) has been used as a technique to directly align LLMs with human preferences (Ziegler et al., 2019) through training a reward model on pairwise comparisons of natural language responses. It has been successfully applied for summarization (Stiennon et al., 2020), generalized instruction following (Ouyang et al., 2022; Lai et al., 2023), dialogue (Gilardi et al., 2023; Manyika, 2023; Glaese et al., 2022; Bai et al., 2022a) and question answering (Nakano et al., 2021).
LLMs have also been extensively used for data generation (Wang et al., 2021b; Meng et al., 2023), augmentation (Feng et al., 2021) and in self-training setups (Wang et al., 2022b; Madaan et al., 2023). Bai et al. (2022b) introduced the idea of RLAIF, which used LLM labeled preferences in conjunction with human labeled preferences to jointly optimize for the two objectives of helpfulness and harmlessness. Recent works have also explored related techniques for generating rewards from LLMs (Roit et al., 2023; Kwon et al., 2022; Yang et al., 2023). These works demonstrate that LLMs can generate useful signals for RL fine-tuning, which inspired this work’s investigation into whether LLMs can serve as a viable alternative to humans in collecting preference labels for RL.
@@ 7 Conclusion
In this work, we show that RLAIF achieves comparable improvements to RLHF on three text generation tasks. Our experiments show that RLAIF greatly improves upon a SFT baseline, and the margin of improvement is on par with or greater than that of RLHF. Furthermore, in head-to-head comparisons, RLAIF and RLHF are preferred at similar rates by humans. Additionally, we show that RLAIF is effective even when the LLM labeler is the same size as the policy, and directly prompting the LLM labeler to provide rewards during RL can outperform the canonical RLAIF setup that distills preferences into a separate RM. Finally, we study the impact of AI labeling techniques on alignment to human preferences.
While this work highlights the potential of RLAIF, there remain many fascinating open questions, such as whether conducting RLAIF iteratively can achieve additional gains (i.e. use the most recent RLAIF policy to generate new response pairs, conduct RLAIF, and repeat), how RLAIF can be adapted to a model-based RL setting where both human and assistant are modeled by LLMs, and how AI feedback can be leveraged for more specific credit assignment. We leave these questions for future work.
Ethics
One ethical consideration concerns the utilization of AI-generated feedback as a source for model alignment. There exists a potential risk of transferring biases from the off-the-shelf LLM into the generated preferences. This in turn may result in RL-trained policies further amplifying biases, thereby inadvertently misaligning models and potentially causing harm. Extreme caution must be exercised, especially when deploying these models in high-stakes domains such as medicine, law, and employment, where they have the potential to significantly impact human lives in adverse ways. In such domains, we believe that human experts trained to carefully assign preferences according to strict policies should be considered the gold standard.
Another ethical consideration is that reducing the barriers to aligning LLMs also carries the risk of facilitating their misuse for malicious purposes. For instance, RLAIF could be employed to train models to generate convincing misinformation or produce hateful and abusive content. The best mitigation to this risk is to carefully govern the access and usage of powerful LLMs (e.g. limiting “white-box” access), to prevent bad actors from misusing them.
Reproducibility
To promote the reproducibility of this work, many of the details of this research are shared throughout the paper. Open-source datasets are elaborated upon in Appendix C, LLM labeling details in Appendix D, the RL formulation in Appendix F, model training details in Appendix G, human evaluation details in I, and the most critical prompts used in the Appendix (e.g. Tables 17, 21, and 22). Please reach out to authors for any additional questions or requests.
PaLM 2 models are available through Google Cloud’s Vertex API, and the experiments in this work may also be repeated with other publicly available LLMs.

### v3

@@ 3.4 Human Evaluation
To measure win rates, evaluators were presented with an input context and multiple responses generated from different policies (e.g. RLAIF, RLHF, and SFT). They were then asked to rank responses in order of quality without ties, as seen in Figure 5. Input contexts were drawn from the test splits of each dataset, which were not used for training or any other evaluation [^5] For summarization, we used the test split of Reddit TL;DR. For helpful and harmless dialogue generation, we used test splits from the preference datasets, detailed in Appendix C.. Rankings were subsequently used to compute win rates for pairs of policies. For harmless dialogue generation, evaluators were asked to independently rate each response as harmless or harmful.
@@ 4 Results
Table 1: Left side: Win rates for pairs of policies on the summarization and the helpful dialogue tasks. Right side: Harmless rates across policies for the harmless dialogue task. All numbers are based on human evaluation.
@@
d-RLAIF vs SFT 74% 66%
d-RLAIF vs Same-size RLAIF 60%
@@
RLAIF achieves performance gains on par with or better than RLHF on all three tasks (see Figure 1 and Table 1). Specifically, RLAIF and RLHF are preferred by human evaluators over the baseline SFT policy 71% and 73% of the time for summarization [^6] RLAIF and RLHF are also preferred over the human reference summaries in Reddit TL;DR 79% and 80% of the time, respectively. and 63% and 64% for helpful dialogue generation, respectively. The difference in win rates between RLAIF vs. SFT and RLHF vs. SFT are not statistically significant. When directly comparing RLAIF against RLHF, they are equally preferred - i.e. the win rate is not statistically significantly different from 50%. For harmless dialogue generation, RLAIF achieves a harmless rate of 88%, outperforming both RLHF and SFT, which score 76% and 64%, respectively [^7] RLAIF achieves a statistically significant improvement over RLHF and SFT, according to two-sided paired t-tests..
@@
Similar to Stiennon et al. (2020), we observe that RLAIF and RLHF policies tend to generate longer responses than the SFT policy, which may bias human evaluation. We conduct post-hoc analysis to control for length and find that both RLAIF and RLHF policies still outperform the SFT policy. See Appendix J for details.
@@
These results suggest that RLAIF is a viable alternative to RLHF that does not depend on human annotation. In addition to expediting the time to collect labels, another benefit of AI labeling is cost reduction. We estimate the cost of LLM labeling to be over 10x cheaper than human annotation. See Appendix L for detailed analysis.
@@ 4.2 Towards Self-Improvement
In Section 4.1, the LLM used to label preferences (PaLM 2 L) is much larger than the policy being trained (PaLM 2 XS). Going one step further, we explore whether RLAIF can yield improvements when the AI labeler is the same size as the policy. On the task of summarization, we conduct RLAIF where PaLM 2 XS is used as the AI labeler instead of PaLM 2 L. The rest of the setup mimics the experiment in Section 4.1. We refer to this setup as “same-size RLAIF”.
Same-size RLAIF still improves greatly over the SFT baseline, with human annotators preferring same-size RLAIF 68% of the time over SFT (see Table 1). For reference, RLAIF using an AI labeler larger than the policy is preferred 71% over SFT [^8] The difference between win rates between “same-size RLAIF vs. SFT” and the original “RLAIF vs. SFT” is not statistically significant. For a two-sample t-test, p-value = 0.07. At alpha = 0.05, this difference is not statistically significant.. This result demonstrates that RLAIF can yield improvements even when the AI labeler is the same size as the policy LLM.
We note that this experiment is not a strict example of “self-improvement” (Huang et al., 2022), as the AI labeler is the instruction-tuned PaLM 2 XS, whereas the initial policy is PaLM 2 XS fine-tuned on Reddit TL;DR summarization. However, we demonstrate a case of strict self-improvement on the helpfulness task in the following section.
4.3 D-RLAIF
In Sections 4.1 and 4.2, AI feedback was distilled into a RM. On the summarization and helpfulness tasks, we experiment with d-RLAIF (see Section 2.2.2). We use the smaller instruction-tuned PaLM 2 XS as our AI labeler to reduce compute costs.
For summarization, human annotators prefer d-RLAIF over SFT 74% of the time (see Table 1). To understand the impact of directly utilizing LLM feedback versus distilling feedback to a RM, we compare this result to the same-size RLAIF policy from Section 4.2, which only differs in terms of the reward function. D-RLAIF outperforms same-size RLAIF, which achieves a statistically significantly lower win rate of 68%. Furthermore, when shown responses side-by-side, annotators prefer d-RLAIF over same-size RLAIF 60% of the time [^9] This is statistically significantly different from 50% according to the binomial test.. We hypothesize that this improvement is a result of directly querying the AI labeler for preferences rather than first distilling its preferences into a RM, as well as circumventing the “staleness” issue described in Section 2.2.2.
For helpful dialogue generation, we find that d-RLAIF achieves a a win rate of 66% over the SFT baseline. Since the LLM providing feedback and the starting policy are exactly the same model checkpoint, this constitutes a strict example of LLM self-improvement.
@@ 4.4 Prompting Techniques
Table 2: We observe that eliciting chain-of-thought reasoning tends to improve AI labeler alignment, while few-shot prompting and detailed preambles have mixed effects across tasks. Above, “Help.” and “Harm.” refer to helpfulness and to harmlessness, respectively.
@@
Prompt Summary Help. Harm.
Base 0-shot 76.1% 67.8% 69.4%
Base 1-shot 76.0% 67.1% 71.7%
Base 2-shot 75.7% 66.8% 72.1%
Base + CoT 0-shot 77.5% 69.1% 70.6%
Detailed 0-shot 77.4% 67.6% 70.1%
Detailed 1-shot 76.2% 67.6% 71.5%
Detailed 2-shot 76.3% 67.3% 71.6%
Detailed 8-shot 69.8%
Detailed + CoT 0-shot 78.0% 67.8% 70.1%
Detailed + CoT 1-shot 77.4% 67.4% 69.9%
Detailed + CoT 2-shot 76.8% 67.4% 69.2%
@@
We experiment with three types of prompting variations - preamble specificity, chain-of-thought reasoning, and in-context learning (see Table 2). The best prompts outperform the base prompts (“Base 0-shot”) by +1.9%, +1.3%, and +1.7% for summarization, helpfulness, and harmlessness, respectively.
@@
Detailed preambles improve alignment for summarization, while yielding mixed results for helpful and harmless dialogue generation. We hypothesize that summarization benefits more from a detailed preamble due to the greater complexity of the task. Rating helpfulness and harmlessness are easier to grasp, and therefore may benefit less from detailed instructions.
@@
Chain-of-thought reasoning generally improves alignment. For summarization, the improvement is consistent. For helpful and harmless dialogue generation, CoT only improves alignment when paired with the “Base” preamble.
@@
Surprisingly, we observe that in-context learning only improves alignment for harmless dialogue generation [^10] We verified that all inputs used in these experiments fit within our AI labeler’s context length.. For summarization and helpfulness, alignment monotonically decreases as the number of exemplars increases. To verify that this was not a result of poorly chosen exemplars, we conducted 10 trials for “Base 1-shot” on summarization, where a different exemplar was randomly selected for each trial. The maximum AI labeler alignment from all trials was 76.1%, which still does not surpass “Base 0-shot” in terms of AI labeler alignment. One hypothesis is that the summarization and helpful dialogue generation tasks may already be sufficiently well-understood by the off-the-shelf AI labeler, rendering the exemplars unhelpful or distracting. It is worth noting that in-context learning is still an important research area that is not fully understood (Min et al., 2022; Wang et al., 2022a).
@@
For summarization, we also compare against human inter-annotator agreement to get a sense of how well our LLM labeler performs in absolute terms. Stiennon et al. (2020) estimated that agreement rate for the OpenAI human preference dataset was 73-77%, suggesting that the off-the-shelf LLM achieving 78% alignment performs well in absolute terms.
@@
We also conduct experiments with self-consistency (Wang et al., 2022b), where multiple chain-of-thought rationales are sampled with temperature $T>0$. The preference distributions generated by the LLM are averaged together to arrive at the final preference label. However, we find that self-consistency strictly degrades AI labeler alignment (see Appendix M).
@@
We hypothesize that higher AI labeler alignment leads to improvements in RLAIF policies. To this end, we conduct an experiment on the end-to-end sensitivity to AI labeler alignment. Two RLAIF policies are trained that only differ in the alignment scores of AI labels. Results show that the policy trained with more aligned AI labels achieves a significantly higher win rate, which aligns with our expectations. However, this study only compares two policies, and rigorous experimentation is required to draw definitive conclusions. See Appendix N for details.
@@ 4.5 Size of LLM Labeler
Model Size AI Labeler Alignment
@@
PaLM 2 S 73.8%
PaLM 2 XS 62.7%
Large model sizes are not widely accessible and can be expensive to run. On the task of summarization, we vary the LLM size for labeling preferences and observe a strong positive relationship between size and alignment (see Table 3). Alignment decreases by 4% when substituting PaLM 2 L with PaLM 2 S, and decreases another 11% when using PaLM 2 XS - a trend consistent with scaling behaviors observed in other work (Kaplan et al., 2020). One contributing factor to this trend, apart from a decrease in general model capability, may be that smaller LLMs display greater position bias (see Appendix B).
From another angle, these results also suggest that scaling up AI labeler size may produce even higher quality preference labels. Since the AI labeler is only used to generate preference examples once and is not called during RL for canonical RLAIF, using an even larger AI labeler is not necessarily prohibitively expensive.
@@ 5 Qualitative Observations
To gain a deeper understanding of how RLAIF compares to RLHF, we visually inspected responses generated by both policies for the summarization task. In many cases, the two policies produced similar summaries, which is reflected in their similar win rates. However, we identified a few patterns where they sometimes diverged.
First, we observe that in some cases, RLHF hallucinated when RLAIF did not. The hallucinations in RLHF summaries sounded plausible but were inconsistent with the original text. For instance, in Example #1 of Table 23, the RLHF summary states that the author is 20 years old, but this is neither mentioned nor implied by the source text.
Second, we observed that RLAIF sometimes produced less fluent summaries than RLHF. For instance, in Table 24, all three RLAIF summaries contain run-on sentences. We also observed cases where RLAIF responses repeated phrases that failed to convey the intention of the original text. For example, several summaries concluded with “How do I get over this?”, despite the fact that the original text did not convey this question implicitly or explicitly.
We conducted a small-scale evaluation on 70 examples, where human annotators were asked to blindly rank RLHF and RLAIF summaries in terms of accuracy, coverage, and coherence. However, the difference in scores was not statistically significant. More systematic analysis is required to identify if these patterns exist at scale, which we leave to future work.
@@ 6 Related Work
LLMs have shown impressive performance on a wide range of NLP tasks (Brown et al., 2020; Thoppilan et al., 2022; Chowdhery et al., 2022; Google et al., 2023; OpenAI, 2023a). For several of these tasks, RL has emerged as an effective optimization technique. While initial applications of RL on tasks such as translation (Wu et al., 2016, 2018) and summarization (Gao et al., 2019; Wu & Hu, 2018) used automatic evaluation metrics as rewards, such simplified formulations of rewards did not fully align with human notions of quality.
Reinforcement learning (RL) from human feedback (Christiano et al., 2017) has been used as a technique to directly align LLMs with human preferences (Ziegler et al., 2019) through training a reward model on pairwise comparisons of natural language responses. It has been successfully applied for summarization (Stiennon et al., 2020), instruction following (Ouyang et al., 2022; Lai et al., 2023), dialogue (Gilardi et al., 2023; Manyika, 2023; Glaese et al., 2022; Bai et al., 2022a) and question answering (Nakano et al., 2021). To mitigate some stability and efficiency challenges of conducting RL, DPO (Rafailov et al., 2024) reformulates the training objective to rely on a classification loss, while RaFT (Dong et al., 2023) employs the reward model for conducting rejection-sampling fine-tuning.
LLMs have also been extensively used for data generation (Wang et al., 2021b; Meng et al., 2023), augmentation (Feng et al., 2021) and in self-training setups (Wang et al., 2022b; Madaan et al., 2023). Bai et al. (2022b) introduced the idea of RLAIF, which used LLM and human labeled preferences together to jointly optimize for the two objectives of helpfulness and harmlessness. Recent works have also explored related techniques for generating rewards from LLMs (Roit et al., 2023; Kwon et al., 2022; Yang et al., 2023). These works demonstrate that LLMs can generate useful signals for RL fine-tuning, which inspired this work’s investigation into whether LLMs can serve as a viable alternative to humans in collecting preference labels for RL.
@@ 7 Conclusion
We show that RLAIF achieves comparable improvements to RLHF on three text generation tasks. In head-to-head comparisons, RLAIF and RLHF are preferred at similar rates by humans. Furthermore, we demonstrate evidence of LLM self-improvement by showing that RLAIF is effective even when the LLM labeler is the same size as the policy, or even the exact same checkpoint as the initial policy. Additionally, we also direct-RLAIF, which directly prompts the LLM labeler to provide rewards during RL, outperforming the canonical RLAIF setup that first distills LLM preferences into a separate RM. Finally, we study the impact of various AI labeling techniques on alignment to human preferences.
While this work highlights the potential of RLAIF, there remain many fascinating open questions, such as how RLAIF can be adapted to a model-based RL setting where both human and assistant are modeled by LLMs, or how AI feedback can be leveraged for granular credit assignment. We leave these questions to future work.